OpenStudy (anonymous):

Find x + y: -3x + 7y = -16 -9x + 5y = 16

OpenStudy (jackellyn):

Solve for y in the first eqn y=(-16 +3x)/7 Substitute into second eqn -9x+5((-16+3x)/7)=16 Distribute the 5 -9x - 80/7 + 15x/7 = 16 Multiple everything by 7 (to get rid of fraction) -9(7)x -80 + 15x = 112 Simplify -63x - 80 +15x = 112 -48x -80 =112 -48x=192 x=-4 Substitute into first eqn -3(-4)+7y=-16 12 + 7y = -16 7y= -28 y = -4 x+y= -4 + -4 = -8